3. A consumer spend all his income on Coke and Biscuits. At the current prices of Px =Rs.100 and Py =Rs.50, he maximizes his utility by purchasing 20 units of Coke and 50 Units of 

Biscuits 

a) What is consumer’s Income? 

b) What is consumer’s marginal rate of substitution of Coke for Biscuits at the equilibrium position? 



1
Expert's answer
2021-02-01T07:48:04-0500

Solution


a) What is consumer’s Income?


Given that the current prices of Coke (Px =Rs.100) and Biscuits (Py =Rs.50), the consumer purchases 20 units of Coke and 50 Units of Biscuits


Consumer's income = Maximum Utility = Rs.(100 x 20) + Rs. (50 x 50) = Rs. 4500


b) What is consumer’s marginal rate of substitution of Coke for Biscuits at the equilibrium position?


MRS along an indifference curve: How much Y is the consumer willing to give up in order to get 1 more of X. Usually shown positive (numerical value)


An important thing is to derive MRS.


"MRS = - \frac{dy}{dx}=" |Slope of Indifference Curve|



"MRS = -\frac{50}{100}=-\frac12"